In two years time I will be 3 times as old as I am now? How old am I?
How do you work this out? and what is the answer?

Answer by fcabanski(1391)

For word problems like this one assign a variable to the unknown (in this case current age) then translate words into math. Is or will be mean =. Three times something means 3 * x. In two years time means plus 2.


You don't know how old you are now, so call it x.


In two years time you are x+2...2 added to your current age.


The problem gives a second relationship between your current age (x) and your age in two years (x+2) ---> you will be 3 times as old as you are now so x+2 = 3*x.


Solve for x: x+2 =3x (subtract x from both sides) 2 = 2x (divide both sides by 2) 1=x. You are 1 years old. So stop worrying about homework. 1 year olds don't have math homework.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)